# Mapping the sources of CP violation in neutrino oscillations from the seesaw mechanism

### v1

@@ 2 Calculations of ${\cal J}_{\nu}$
where $c_{ij}\equiv\cos\theta_{ij}$, $s_{ij}\equiv\sin\theta_{ij}$ and $\hat{s}_{ij}\equiv s_{ij}e^{{\rm i}\delta_{ij}}$ with $\theta_{ij}$ lying in the first quadrant. Some constraints on the 9 active-sterile flavor mixing angles $\theta_{ij}$ (for $i=1,2,3$ and $j=4,5,6$) can be obtained from the currently available bounds on non-unitarity of the PMNS lepton flavor mixing matrix $U$ [22, 23, 24, 25] via the relations $AA^{\dagger}=UU^{\dagger}$ and $RR^{\dagger}=I-AA^{\dagger}$. A conservative expectation is that the magnitudes of $\theta_{ij}$ are at least smaller than ${\cal O}(10^{-1.5})$, assuring that $A\simeq I$ and $U\simeq U_{0}$ hold to a good degree of accuracy, together with
$A^{-1}R\simeq R\simeq\left(\begin{matrix}\hat{s}^{*}_{14}&\hat{s}^{*}_{15}&\hat{s}^{*}_{16}\cr\hat{s}^{*}_{24}&\hat{s}^{*}_{25}&\hat{s}^{*}_{26}\cr\hat{s}^{*}_{34}&\hat{s}^{*}_{35}&\hat{s}^{*}_{36}\cr\end{matrix}\right)+{\cal O}(s^{3}_{ij})\;,$ (6)
where $s_{ij}\equiv\sin\theta_{ij}$ and $\hat{s}_{ij}\equiv s_{ij}e^{{\rm i}\delta_{ij}}$ with $\theta_{ij}$ lying in the first quadrant. A proper redefinition of the phases of three charged lepton fields in Eq. (1) allows one to rotate away 3 of the 9 phase parameters (or their independent combinations) of $A$ and $R$. We are therefore left with 6 nontrivial CP-violating phases which can be chosen from
@@
where $M_{\nu}\equiv U_{0}D_{\nu}U^{T}_{0}$ and $\Delta_{ii^{\prime}}\equiv m^{2}_{i}-m^{2}_{i^{\prime}}$ (for $i,i^{\prime}=1,2,3$). One may derive the expressions of $m^{2}_{i}$ by use of the parameters of $D_{N}$ and $A^{-1}R$ from Eq. (2), and then arrive at those of $\Delta_{ii^{\prime}}$. But a global analysis of the neutrino oscillation data has given the best-fit values $\Delta_{21}\simeq 7.41\times 10^{-5}~{}{\rm eV}^{2}$ and $\Delta_{31}\simeq 2.51\times 10^{-3}~{}{\rm eV}^{2}$ (normal mass ordering) or $\Delta_{31}\simeq-2.41\times 10^{-3}~{}{\rm eV}^{2}$ (inverted mass ordering) [31, 32]. So the product $\Delta_{21}\Delta_{31}\Delta_{32}$ on the right-hand side of Eq. (8) is already known to us, and it is insensitive to the sign ambiguities of $\Delta_{31}$ and $\Delta_{32}$. In this case what we need to do is to calculate $(M_{\nu}M^{\dagger}_{\nu})_{e\mu}$, $(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}$ and $(M_{\nu}M^{\dagger}_{\nu})_{\tau e}$ in terms of the original seesaw parameters $M_{j}$, $\theta_{ij}$ and $\delta_{ij}$ (for $i=1,2,3$ and $j=4,5,6$) with the help of Eqs. (2) and (6):
@@
(1) For the sake of simplicity, we denote a sum of the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ proportional to $M^{3}_{j}M^{3}_{k}$ (for $k\neq j$) as $T_{\bf 33}$, and a sum of those terms proportional to $M^{4}_{j}M^{2}_{k}$ (for $k\neq j$) as $T_{\bf 42}$. Our results of $T_{\bf 33}$ and $T_{\bf 42}$ are summarized as
@@
(2) The terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ proportional to $M^{4}_{j}M_{k}M_{l}$ (for $l\neq k\neq j$) are summarily denoted as $T_{\bf 411}$. The explicit analytical result of $T_{\bf 411}$ is found to be
@@
(3) A sum of the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ that are proportional to $M^{3}_{4}M^{2}_{5}M_{6}$, $M^{3}_{4}M^{2}_{6}M_{5}$, $M^{3}_{5}M^{2}_{4}M_{6}$, $M^{3}_{5}M^{2}_{6}M_{4}$, $M^{3}_{6}M^{2}_{4}M_{5}$ and $M^{3}_{6}M^{2}_{5}M_{4}$ is denoted as $T_{\bf 321}$. The explicit analytical result of $T_{\bf 321}$ is obtained as follows:
@@
(4) Finally, the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ that are proportional to $M^{2}_{4}M^{2}_{5}M^{2}_{6}$ are summarily denoted as $T_{\bf 222}$. The explicit analytical result of $T_{\bf 222}$ is found to be
@@
$+\left(\left|I_{45}\right|^{2}I_{66}-\left|I_{64}\right|^{2}I_{55}\right)\Bigg{[}\sum^{3}_{i=1}s^{2}_{i4}s_{i^{\prime}5}s_{i^{\prime}6}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\sin\left(\beta_{i^{\prime}}-\beta_{i^{\prime\prime}}\right)\Bigg{]}$
$+I_{44}\Bigg{[}\sum^{3}_{i=1}s_{i5}s_{i6}\Big{[}s_{i5}s_{i6}\cos 2\beta_{i}+2s_{i^{\prime}5}s_{i^{\prime}6}\cos\left(\beta_{i}+\beta_{i^{\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}6}\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}4}\sin\left(\alpha_{i}-\beta_{i^{\prime}}+\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}4}s_{i^{\prime\prime}5}\sin\left(\alpha_{i}-\beta_{i^{\prime\prime}}+\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}$
$-I_{44}\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}6}\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}4}\cos\left(\alpha_{i}-\beta_{i^{\prime}}+\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}4}s_{i^{\prime\prime}5}\sin\left(\alpha_{i}-\beta_{i^{\prime\prime}}+\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s_{i5}s_{i6}\Big{[}s_{i5}s_{i6}\sin 2\beta_{i}+2s_{i^{\prime}5}s_{i^{\prime}6}\sin\left(\beta_{i}+\beta_{i^{\prime}}\right)\Big{]}\Bigg{]}$
@@
where the analytical results of $T_{\bf 33}$, $T_{\bf 42}$, $T_{\bf 411}$, $T_{\bf 321}$ and $T_{\bf 222}$ have been given in Eqs. (11)—(15). Let us emphasize that these results are completely new. It is certainly too lengthy to write out all the abbreviation terms in the above equations, but one may figure out all the combinations of the CP-violating phases $\alpha_{i}$, $\beta_{i}$ and $\gamma_{i}$ (for $i=1,2,3$) that appear in the expression of ${\cal J}_{\nu}$. After a very tedious survey of all the possibilities, we find that there are totally 240 linear combinations of the 6 original seesaw phase parameters:
@@
if the third heavy Majorana neutrino species $N_{6}$ is switched off, in which case the canonical seesaw mechanism is simplified to the so-called minimal seesaw scenario [34, 35, 36, 37]. Such an amazing simplification of the expression of ${\cal J}_{\nu}$ implies that a lot of information about CP violation in the canonical case is lost in the minimal seesaw framework, although the latter is a good benchmark model.

### v2

@@ 2 Calculations of ${\cal J}_{\nu}$
where $c_{ij}\equiv\cos\theta_{ij}$, $s_{ij}\equiv\sin\theta_{ij}$ and $\hat{s}_{ij}\equiv s_{ij}e^{{\rm i}\delta_{ij}}$ with $\theta_{ij}$ lying in the first quadrant. As a consequence,
$A^{-1}R=\left(\begin{matrix}\hat{t}^{*}_{14}&c^{-1}_{14}\hat{t}^{*}_{15}&c^{-1}_{14}c^{-1}_{15}\hat{t}^{*}_{16}\cr\vspace{-0.45cm}\cr c^{-1}_{14}\hat{t}^{*}_{24}&\hat{t}_{14}\hat{t}^{*}_{15}\hat{t}^{*}_{24}+c^{-1}_{15}c^{-1}_{24}\hat{t}^{*}_{25}&\begin{array}[]{l}+\hat{t}_{14}c^{-1}_{15}\hat{t}^{*}_{16}\hat{t}^{*}_{24}+\hat{t}_{15}\hat{t}^{*}_{16}c^{-1}_{24}\hat{t}^{*}_{25}\\ +c^{-1}_{16}c^{-1}_{24}c^{-1}_{25}\hat{t}^{*}_{26}\end{array}\cr\vspace{-0.45cm}\cr c^{-1}_{14}c^{-1}_{24}\hat{t}^{*}_{34}&\begin{array}[]{l}+\hat{t}_{14}\hat{t}^{*}_{15}c^{-1}_{24}\hat{t}^{*}_{34}+c^{-1}_{15}\hat{t}_{24}\hat{t}^{*}_{25}\hat{t}^{*}_{34}\\ +c^{-1}_{15}c^{-1}_{25}c^{-1}_{34}\hat{t}^{*}_{35}\end{array}&\begin{array}[]{l}+\hat{t}_{14}c^{-1}_{15}\hat{t}^{*}_{16}c^{-1}_{24}\hat{t}^{*}_{34}+\hat{t}_{15}\hat{t}^{*}_{16}\hat{t}_{24}\hat{t}^{*}_{25}\hat{t}^{*}_{34}\\ +\hat{t}_{15}\hat{t}^{*}_{16}c^{-1}_{25}c^{-1}_{34}\hat{t}^{*}_{35}+c^{-1}_{16}\hat{t}_{24}c^{-1}_{25}\hat{t}^{*}_{26}\hat{t}^{*}_{34}\\ +c^{-1}_{16}\hat{t}_{25}\hat{t}^{*}_{26}c^{-1}_{34}\hat{t}^{*}_{35}+c^{-1}_{16}c^{-1}_{26}c^{-1}_{34}c^{-1}_{35}\hat{t}^{*}_{36}\end{array}\cr\end{matrix}\right)\;,$ (6)
where $\hat{t}_{ij}\equiv e^{{\rm i}\delta_{ij}}\tan\theta_{ij}$ is defined. Some constraints on the 9 active-sterile flavor mixing angles $\theta_{ij}$ (for $i=1,2,3$ and $j=4,5,6$) have been obtained from the currently available bounds on non-unitarity of the PMNS lepton flavor mixing matrix $U$ [22, 23, 24, 25] via the relations $AA^{\dagger}=UU^{\dagger}$ and $RR^{\dagger}=I-AA^{\dagger}$, from which one finds that the magnitudes of $\theta_{ij}$ must be smaller than ${\cal O}(10^{-1.5})$. So a Maclaurin expansion of Eq. (6) leads us to
$A^{-1}R=\left(\begin{matrix}\hat{s}^{*}_{14}&\hat{s}^{*}_{15}&\hat{s}^{*}_{16}\cr\hat{s}^{*}_{24}&\hat{s}^{*}_{25}&\hat{s}^{*}_{26}\cr\hat{s}^{*}_{34}&\hat{s}^{*}_{35}&\hat{s}^{*}_{36}\cr\end{matrix}\right)+{\cal O}(s^{3}_{ij})+\cdots\;.$ (7)
Keeping only the leading term of $A^{-1}R$ turns out to be a very good approximation, independent of any possible hierarchy or degeneracy of the original seesaw parameters. The smallness of active-sterile flavor mixing implies that the general Jarlskog invariants of CP violation defined in terms of $U$ is actually dominated by the unique ${\cal J}_{\nu}$ defined in Eq. (3) with the help of $U_{0}$, namely
${\cal J}^{ii^{\prime}}_{\alpha\beta}\equiv{\rm Im}\left(U_{\alpha i}U_{\beta i^{\prime}}U^{*}_{\alpha i^{\prime}}U^{*}_{\beta i}\right)={\cal J}_{\nu}+\Delta{\cal J}^{ii^{\prime}}_{\alpha\beta}\;,$ (8)
where $(\alpha,\beta)$ run cyclically over $(e,\mu,\tau)$, $(i,i^{\prime})$ run cyclically over $(1,2,3)$, and $\Delta{\cal J}^{ii^{\prime}}_{\alpha\beta}$ stand for the corresponding small corrections to ${\cal J}_{\nu}$. Taking account of the recent numerical bounds on possible deviations of $AA^{\dagger}$ from $I$ [25] and the fact that $\Delta{\cal J}^{ii^{\prime}}_{\alpha\beta}$ are at least suppressed by the factors of ${\cal O}(s^{2}_{ij})$ (for $i=1,2,3$ and $j=4,5,6$[12], we find that $|\Delta{\cal J}^{ii^{\prime}}_{\alpha\beta}/{\cal J}_{\nu}|\lesssim 10^{-2}$ is a safe estimate especially for $|{\cal J}_{\nu}|\sim 10^{-2}$ as favored by the present T2K data [26, 27]. In other words, ${\cal J}^{ii^{\prime}}_{\alpha\beta}\simeq{\cal J}_{\nu}$ is also expected to be a very good approximation. So it will make sense to substitute the approximate expression of $A^{-1}R$ in Eq. (7) into the seesaw formula in Eq. (2) to calculate all the leading terms of ${\cal J}_{\nu}$.
A proper redefinition of the phases of three charged lepton fields in Eq. (1) allows one to rotate away 3 of the 9 phase parameters (or their independent combinations) of $A$ and $R$. We are therefore left with 6 nontrivial CP-violating phases which can be chosen from
@@
where $M_{\nu}\equiv U_{0}D_{\nu}U^{T}_{0}$ and $\Delta_{ii^{\prime}}\equiv m^{2}_{i}-m^{2}_{i^{\prime}}$ (for $i,i^{\prime}=1,2,3$). One may derive the expressions of $m^{2}_{i}$ by use of the parameters of $D_{N}$ and $A^{-1}R$ from Eq. (2), and then arrive at those of $\Delta_{ii^{\prime}}$. But a global analysis of the neutrino oscillation data has given the best-fit values $\Delta_{21}\simeq 7.41\times 10^{-5}~{}{\rm eV}^{2}$ and $\Delta_{31}\simeq 2.51\times 10^{-3}~{}{\rm eV}^{2}$ (normal mass ordering) or $\Delta_{31}\simeq-2.41\times 10^{-3}~{}{\rm eV}^{2}$ (inverted mass ordering) [31, 32]. So the product $\Delta_{21}\Delta_{31}\Delta_{32}$ on the right-hand side of Eq. (10) is already known to us, and it is insensitive to the sign ambiguities of $\Delta_{31}$ and $\Delta_{32}$. In this case what we need to do is to calculate $(M_{\nu}M^{\dagger}_{\nu})_{e\mu}$, $(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}$ and $(M_{\nu}M^{\dagger}_{\nu})_{\tau e}$ in terms of the original seesaw parameters $M_{j}$, $\theta_{ij}$ and $\delta_{ij}$ (for $i=1,2,3$ and $j=4,5,6$) with the help of Eqs. (2) and (7):  [^3] If the exact expression of $A^{-1}R$ in Eq. (6) is used to calculate ${\cal J}_{\nu}$ with the help of Eqs. (2) and (10), much more independent phase combinations are expected to appear in the sub-leading terms of the analytical expression of ${\cal J}_{\nu}$. But such terms are suppressed by the higher-order factors of the small active-sterile flavor mixing angles, as can be easily seen from Eq. (7), and hence their effects are negligibly small.
@@
2.1 Terms $T_{\bf 33}$ and $T_{\bf 42}$
For the sake of simplicity, we denote a sum of the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ proportional to $M^{3}_{j}M^{3}_{k}$ (for $k\neq j$) as $T_{\bf 33}$, and a sum of those terms proportional to $M^{4}_{j}M^{2}_{k}$ (for $k\neq j$) as $T_{\bf 42}$. Our results of $T_{\bf 33}$ and $T_{\bf 42}$ are summarized as
@@ 2.1 Terms $T_{\bf 33}$ and $T_{\bf 42}$
2.2 Term $T_{\bf 411}$
The terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ proportional to $M^{4}_{j}M_{k}M_{l}$ (for $l\neq k\neq j$) are summarily denoted as $T_{\bf 411}$. The explicit analytical result of $T_{\bf 411}$ is found to be
@@ 2.2 Term $T_{\bf 411}$
2.3 Term $T_{\bf 321}$
Let us denote a sum of the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ proportional to $M^{3}_{4}M^{2}_{5}M_{6}$, $M^{3}_{4}M^{2}_{6}M_{5}$, $M^{3}_{5}M^{2}_{4}M_{6}$, $M^{3}_{5}M^{2}_{6}M_{4}$, $M^{3}_{6}M^{2}_{4}M_{5}$ and $M^{3}_{6}M^{2}_{5}M_{4}$ as $T_{\bf 321}$. The explicit analytical result of $T_{\bf 321}$ is obtained as follows:
@@ 2.3 Term $T_{\bf 321}$
2.4 Term $T_{\bf 222}$
Finally, the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ that are proportional to $M^{2}_{4}M^{2}_{5}M^{2}_{6}$ are summarily denoted as $T_{\bf 222}$. The explicit analytical result of $T_{\bf 222}$ is found to be
@@ 2.4 Term $T_{\bf 222}$
$+\left(I_{55}\left|I_{64}\right|^{2}-I_{44}\left|I_{56}\right|^{2}\right)\Bigg{[}\sum^{3}_{i=1}s^{2}_{i6}s_{i^{\prime}4}s_{i^{\prime}5}s_{i^{\prime\prime}4}s_{i^{\prime\prime}5}\sin\left(\alpha_{i^{\prime}}-\alpha_{i^{\prime\prime}}\right)\Bigg{]}$
$-I_{66}\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}\Big{[}s_{i4}s_{i5}\cos 2\alpha_{i}+2s_{i^{\prime}4}s_{i^{\prime}5}\cos\left(\alpha_{i}+\alpha_{i^{\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}6}\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}4}\sin\left(\alpha_{i}-\beta_{i^{\prime}}-\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}4}s_{i^{\prime\prime}5}\sin\left(\alpha_{i}-\beta_{i^{\prime\prime}}-\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}$
$-I_{66}\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}6}\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}4}\cos\left(\alpha_{i}-\beta_{i^{\prime}}-\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}4}s_{i^{\prime\prime}5}\cos\left(\alpha_{i}-\beta_{i^{\prime\prime}}-\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}\Big{[}s_{i4}s_{i5}\sin 2\alpha_{i}+2s_{i^{\prime}4}s_{i^{\prime}5}\sin\left(\alpha_{i}+\alpha_{i^{\prime}}\right)\Big{]}\Bigg{]}$
@@
where the analytical results of $T_{\bf 33}$, $T_{\bf 42}$, $T_{\bf 411}$, $T_{\bf 321}$ and $T_{\bf 222}$ have been given in Eqs. (13)—(17).
Let us emphasize that the analytical results obtained above are completely new, and they provide the first general and explicit expression of ${\cal J}_{\nu}$ in terms of the original flavor parameters in the canonical seesaw mechanism. It is certainly too lengthy to write out all the abbreviation terms in the above equations, but one may figure out all the combinations of the CP-violating phases $\alpha_{i}$, $\beta_{i}$ and $\gamma_{i}$ (for $i=1,2,3$) that appear in the expression of ${\cal J}_{\nu}$.
2.5 Phase combinations
After a very tedious survey of all the possibilities, including a careful count of the phase combinations in those abbreviation terms, we find that there are totally 240 linear combinations of the 6 original seesaw phase parameters in ${\cal J}_{\nu}$:
@@ 2.5 Phase combinations
if the third heavy Majorana neutrino species $N_{6}$ is switched off, in which case the canonical seesaw mechanism is simplified to the so-called minimal seesaw scenario [34, 35, 36, 37]  [^4] In Ref. [38] the authors have successfully derived the exact correlation between the Jarlskog invariant and the CP-violating asymmetries of heavy Majorana neutrino decays by using a different parametrization of the flavor structure for the minimal seesaw mechanism, although their results are not as intuitional as the approximate ones obtained in the Euler-like block parametrization of the minimal seesaw scenario [33]. It is extremely complicated, however, to do the same thing in the canonical seesaw framework containing three heavy Majorana neutrino fields. The present work makes the first attempt in this regard, but a compromised balance has been struck between calculability of the observable quantities and complexity of their analytical expressions (i.e., we have adopted the approximate expressions of $A^{-1}R$ in our calculations from the very beginning). Even in this situation our analytical results have clearly shown how complicated the canonical (3+3) flavor case is, as compared with the minimal (3+2) flavor case.. Such an amazing simplification of the expression of ${\cal J}_{\nu}$ implies that a lot of information about CP violation in the canonical case is lost in the minimal seesaw framework, although the latter is a good benchmark model.
@@ 4 Conclusion
Encouraged by this proof-of-concept work, we are going to calculate all the derivational flavor parameters — three masses, three flavor mixing angles and three CP-violating phases of the light Majorana neutrinos by using the original flavor parameters in the canonical seesaw mechanism [52], such that its parameter space can be constrained by both more precise neutrino oscillation data and more accurate non-oscillation measurements in the near future. Instead of testing many specific seesaw models which are based on various special assumptions, our approach will offer a model-independent way to test the canonical seesaw mechanism as a whole although the number of observable quantities at low energies cannot match the number of the original seesaw flavor parameters. Such an attempt is fundamentally meaningful in the era of precision measurements.